# Few TensoRF: Enhance the Few-shot on Tensorial Radiance Fields

### 26 Mar 2026 v1

@@ 1 Introduction
The ability to create accurate and versatile 3D reconstructions has become increasingly vital in today’s technological landscape. From applications in computer vision to healthcare, and even in the world of entertainment, 3D reconstruction plays a pivotal role. The history of 3D reconstruction methods has witnessed significant advancements, with a continuous pursuit of techniques that can transform single or multiple 2D images into detailed 3D representations. This study, inspired by the success of the renowned Neural Radiance Field (NeRF) mildenhall2020nerf ; brualla2021nerf for synthesizing novel views of complex scenes, aims to incorporate two improved methods related to NeRF: the Tensorial Radiance Field (TensorRF) chen2022tensorf and the enhancement of few-shot neural rendering with Free Frequency Regularization (FreeNeRF) yang2023freenerf . TensoRF introduces a novel approach using 4D tensors and decompositions caroll1970analysis ; lathauwer2008decomposition , for improved rendering quality, reduced memory usage, and faster reconstruction times compared to NeRF. On the other hand, FreeNeRF simplifies few-shot neural rendering by introducing costeffective regularization terms, outperforming complex alternatives with a single line of code change, and emphasizing the significance of frequency in NeRF’s training. This study is motivated by the pressing need for a 3D reconstruction method that is both versatile and efficient. Our primary objective is to develop a versatile 3D reconstruction method, leveraging the combined capabilities of TensorRF and FreeNeRF. This approach is designed to address the shortcomings of previous methods and promote broader applications in various domains, especially experience on human body datasets. In developing our approach, we build upon the foundation of TensoRF and introduce two primary regularization techniques, frequency, and occlusion, inspired by the FreeNeRF paper. These techniques are incorporated as a positional encoding step to map each input 5D coordinate within the tensor component of the scene. Our methodology will be thoroughly compared with previous NeRF-related methods, not only on typical objects but also on the more complex task of 3D human body reconstruction. The THuman2.0 dataset yu2021function4d will be used and trained on different setups for experiments and comparison.
@@ 2.1 Neural Radiance Fields (NeRF)
Neural Radiance Fields (NeRF) mildenhall2020nerf has transformed the approach to 3D reconstruction. Utilizing fully connected deep networks, NeRF processes single continuous 5D coordinates, including spatial location (x, y, z) and viewing direction $(\theta,\phi$) to generate volume density and view-dependent emitted radiance. These properties facilitate the synthesis of novel views by querying 5D coordinates along camera rays and utilizing volume rendering techniques. NeRF’s applications range from novel view synthesis brualla2021nerf , 3D generation poole2022dream ; jain2022zeroshot to deformation park2021nerfies ; pumarola2020dnerf and video li2020neuralscene ; Xian2020spacetime . Despite its remarkable capabilities in high-quality scene representation, NeRF’s significant drawback is its demand for a substantial number of input images, rendering it ineffective at synthesizing novel views from a limited set of input views, such as 3, 6, or 9 views yang2023freenerf . This limitation not only hinders its potential real-world applications but also contributes to its relatively slow rendering speed, with NeRF requiring approximately 35 hours of training for completeness chen2022tensorf .
@@ 2.2 Tensorial Radiance Field (TensorRF)
Tensorial Radiance Fields (TensorRF) chen2022tensorf presents a novel approach to 3D reconstruction by representing the radiance field as a 4D tensor. TensorRF utilizes the tensor decomposition technique Aaneas2016largescale to achieve improved rendering quality, reduced memory usage, and faster reconstruction times compared to NeRF. Although the experience and comparison between the classic CP decomposition kolda2009tensor and a new vector matrix (VM) decomposition, TensorRF focuses mainly on the reconstruction through gradient descent rather than the decomposition itself. This introduces low-rank regularization in the optimization process, ultimately leading to enhanced rendering quality. TensorRF’s advantage lies in its potential to make 3D reconstruction more memory-efficient while maintaining rendering quality.
@@ 2.3 FreeNeRF: Few-shot Neural Rendering
Few-shot neural rendering has posed a challenge, often requiring complex external information or pretraining. FreeNeRF yang2023freenerf addresses this challenge by introducing cost-effective regularization terms, notably frequency and occlusion regularization. Remarkably, these modifications can be incorporated with minimal changes to the original NeRF code, maintaining the same computational efficiency. This approach underscores the importance of geometry in few-shot neural rendering, achieving impressive results without expensive pre-training or complex training-time patch rendering.
@@ 2.4 Human Body 3D Reconstruction
Reconstructing the 3D Human Body is the focus of our latest experiment, aimed at expanding the capabilities of NeRF-like methods beyond their traditional applications on standard datasets like Blender, DTU Aaneas2016largescale ; jensen2014largescale , or LLFF mildenhall2019locallight . Unlike the objects found in these datasets, the challenges posed by a comprehensive Human Body dataset such as Thuman 2.0 yu2021function4d are distinct, ranging from the diverse human shapes and clothing variations to the multitude of poses encountered in real-world scenarios. Our innovative approach, rooted in NeRF-like principles, prioritize generating novel views over constructing the 3D mesh or voxel itself— a departure from earlier methods like ICON xiu2022icon , ECON xiu2023econ and others. This unique perspective holds promise, particularly in scenarios with limited resources, offering a viable solution to address the complexities inherent in capturing the richness of the human form.
@@ 3 Method
The proposed Few-TensoRF is still a 3D reconstruction method, capable of exporting 3D object meshes due to its special 3D point coordinate and RGB with density mapping. However, it functions as just one core process within the broader NeRF application-level pipeline, as depicted in Fig. 1. The Few TensoRF method mainly takes advantage of the strong point in 2 methods: faster training time with TensoRF chen2022tensorf and improvement of the FreeNeRF yang2023freenerf in sparse input cases.
@@ 3.1 Fast training time with Tensorial Radiance Field (TensoRF) Base
Figure 2: The Few TensoRF overview. This illustration was adapted from the original TensoRF paper chen2022tensorf . It highlights two significant enhancements as we move from left to right: incorporating density frequency masks and integrating appearance color frequency on each rendered voxel. Additionally, a supplementary frequency mask is applied to refine the positional encoding fed into the neural networks.
As detailed in the original TensoRF paper chen2022tensorf , the function $S$ (in Eqn. 1) is introduced as a pre-selected function. This function, which can be a small MLP or a spherical harmonics (SH) function, is employed to convert an appearance feature vector and a viewing direction $d$ into the final color $c$. However, in this study, we will take advantage of the variant of MLP structures to get more customization for few-shot scenarios. Those $A$ and $A_{c}$ tensor components play a pivotal role in the revolution of our Few TensorRF technique.
@@ 3.2 Enhance Few-shot on Tensorial Radiance Field (Few TensoRF)
Given the findings depicted in Figure 3, showcasing the rendered novel views with varying amounts of training images in the TensoRF baseline, it becomes evident that TensoRF encounters challenges in delivering satisfactory results under sparse input conditions. The visual comparison highlights a noticeable disparity in the quality of novel views, particularly as the number of training images decreases. As discussed in the FreeNeRF yang2023freenerf , we anticipate encountering issues arising from over-fast convergence during the optimization process. This rapid convergence, particularly in high-frequency components, hinders TensorRF’s ability to effectively explore low frequency information. Consequently, it introduces a quite bias in TensoRF, leading to the manifestation of undesired high-frequency artifacts (TensoRF 3img in Fig. 3).
Therefore, to achieve better results in sparse input cases, the proposed Few TensoRF method proposes three improvements on the TensoRF chen2022tensorf baseline, as depicted in Figure 2. Frequency Masking Tensor Components. The goal here is to decrease the sensitivity of each tensor component in the high-frequency domain during the initial stages of training. This approach guides our model to concentrate on enhancing low-frequency structures, ensuring the stability of tensor components against undesired high-frequency artifacts. As mentioned earlier, there are two types of components: A for the density component and Ac for the appearance color feature component. Two frequency masks will be applied to them according to the following equations:
@@
Where $\alpha_{i}(t,T,L)$ denotes the $i^{th}$ bit value of $\alpha(t,T,L)$ ; $t$ and $T$ are the current training iteration and the total iterations, respectively, $L$ corresponds to either the density component length or the appearance feature length, depending on the desired tensor component to mask. This dynamic frequency mask, a unique feature introduced in FreeNeRF yang2023freenerf , calculates the mask values based on the current iteration if $t$ is less than $T$. The calculation involves both an integer and a fractional part, ensuring a smooth transition. If $t$ is greater than or equal to $T$, the method returns a tensor of ones. This dynamic frequency mask is particularly beneficial for enhancing the high-frequency details of rendering novel views in the later stages of training. Additionally, we have implemented a simplified version of Eqn. 4, allowing us to set a fixed visible ratio for the frequency mask. This can be expressed as follows:
@@ 4.2 Comparison
Table 2: PSNR metrics comparison on each Synthesis NeRF’s scene. The best-performing and second-best results from our experiments are denoted by red and yellow highlights, respectively.

### 30 Mar 2026 v3

@@ 1 Introduction
The ability to create accurate and versatile 3D reconstructions has become increasingly vital in today’s technological landscape. From applications in computer vision to healthcare, and even in the world of entertainment, 3D reconstruction plays a pivotal role. The history of 3D reconstruction methods has witnessed significant advancements, with a continuous pursuit of techniques that can transform single or multiple 2D images into detailed 3D representations. This study, inspired by the success of the renowned Neural Radiance Field (NeRF) Mildenhall et al. (2020); Martin-Brualla et al. (2021) for synthesizing novel views of complex scenes, aims to incorporate two improved methods related to NeRF: the Tensorial Radiance Field (TensorRF) Chen et al. (2022) and the enhancement of few-shot neural rendering with Free Frequency Regularization (FreeNeRF) Yang et al. (2023). TensoRF introduces a novel approach using 4D tensors and decompositions Carroll and Chang (1970); De Lathauwer (2008), for improved rendering quality, reduced memory usage, and faster reconstruction times compared to NeRF. On the other hand, FreeNeRF simplifies few-shot neural rendering by introducing costeffective regularization terms, outperforming complex alternatives with a single line of code change, and emphasizing the significance of frequency in NeRF’s training. This study is motivated by the pressing need for a 3D reconstruction method that is both versatile and efficient. Our primary objective is to develop a versatile 3D reconstruction method, leveraging the combined capabilities of TensorRF and FreeNeRF. This approach is designed to address the shortcomings of previous methods and promote broader applications in various domains, especially experience on human body datasets. In developing our approach, we build upon the foundation of TensoRF and introduce two primary regularization techniques, frequency, and occlusion, inspired by the FreeNeRF paper. These techniques are incorporated as a positional encoding step to map each input 5D coordinate within the tensor component of the scene. Our methodology will be thoroughly compared with previous NeRF-related methods, not only on typical objects but also on the more complex task of 3D human body reconstruction. The THuman2.0 dataset Yu et al. (2021) will be used and trained on different setups for experiments and comparison.
@@ 2.1 Neural Radiance Fields (NeRF)
Neural Radiance Fields (NeRF) Mildenhall et al. (2020) has transformed the approach to 3D reconstruction. Utilizing fully connected deep networks, NeRF processes single continuous 5D coordinates, including spatial location (x, y, z) and viewing direction $(\theta,\phi$) to generate volume density and view-dependent emitted radiance. These properties facilitate the synthesis of novel views by querying 5D coordinates along camera rays and utilizing volume rendering techniques. NeRF’s applications range from novel view synthesis Martin-Brualla et al. (2021), 3D generation Poole et al. (2022); Jain et al. (2022) to deformation Park et al. (2022); Pumarola et al. (2020) and video Li et al. (2020); Xian et al. (2020). Despite its remarkable capabilities in high-quality scene representation, NeRF’s significant drawback is its demand for a substantial number of input images, rendering it ineffective at synthesizing novel views from a limited set of input views, such as 3, 6, or 9 views Yang et al. (2023). This limitation not only hinders its potential real-world applications but also contributes to its relatively slow rendering speed, with NeRF requiring approximately 35 hours of training for completeness Chen et al. (2022).
@@ 2.2 Tensorial Radiance Field (TensorRF)
Tensorial Radiance Fields (TensorRF) Chen et al. (2022) presents a novel approach to 3D reconstruction by representing the radiance field as a 4D tensor. TensorRF utilizes the tensor decomposition technique Aanæs et al. (2016) to achieve improved rendering quality, reduced memory usage, and faster reconstruction times compared to NeRF. Although the experience and comparison between the classic CP decomposition Kolda and Bader (2009) and a new vector matrix (VM) decomposition, TensorRF focuses mainly on the reconstruction through gradient descent rather than the decomposition itself. This introduces low-rank regularization in the optimization process, ultimately leading to enhanced rendering quality. TensorRF’s advantage lies in its potential to make 3D reconstruction more memory-efficient while maintaining rendering quality.
@@ 2.3 FreeNeRF: Few-shot Neural Rendering
Few-shot neural rendering has posed a challenge, often requiring complex external information or pretraining. FreeNeRF Yang et al. (2023) addresses this challenge by introducing cost-effective regularization terms, notably frequency and occlusion regularization. Remarkably, these modifications can be incorporated with minimal changes to the original NeRF code, maintaining the same computational efficiency. This approach underscores the importance of geometry in few-shot neural rendering, achieving impressive results without expensive pre-training or complex training-time patch rendering.
@@ 2.4 Human Body 3D Reconstruction
Reconstructing the 3D Human Body is the focus of our latest experiment, aimed at expanding the capabilities of NeRF-like methods beyond their traditional applications on standard datasets like Blender, DTU Aanæs et al. (2016); Jensen et al. (2014), or LLFF Mildenhall et al. (2019). Unlike the objects found in these datasets, the challenges posed by a comprehensive Human Body dataset such as THuman 2.0 Yu et al. (2021) are distinct, ranging from the diverse human shapes and clothing variations to the multitude of poses encountered in real-world scenarios. Our innovative approach, rooted in NeRF-like principles, prioritize generating novel views over constructing the 3D mesh or voxel itself— a departure from earlier methods like ICON Xiu et al. , ECON Xiu et al. (2023) and others. This unique perspective holds promise, particularly in scenarios with limited resources, offering a viable solution to address the complexities inherent in capturing the richness of the human form.
@@ 3 Method
The proposed Few-TensoRF is still a 3D reconstruction method, capable of exporting 3D object meshes due to its special 3D point coordinate and RGB with density mapping. However, it functions as just one core process within the broader NeRF application-level pipeline, as depicted in Fig. 1. The Few TensoRF method mainly takes advantage of the strong point in 2 methods: faster training time with TensoRF Chen et al. (2022) and improvement of the FreeNeRF Yang et al. (2023) in sparse input cases.
@@ 3.1 Fast training time with Tensorial Radiance Field (TensoRF) Base
Figure 2: The Few TensoRF overview. This illustration was adapted from the original TensoRF paper Chen et al. (2022). It highlights two significant enhancements as we move from left to right: incorporating density frequency masks and integrating appearance color frequency on each rendered voxel. Additionally, a supplementary frequency mask is applied to refine the positional encoding fed into the neural networks.
As detailed in the original TensoRF paper Chen et al. (2022), the function $S$ (in Eqn. 1) is introduced as a pre-selected function. This function, which can be a small MLP or a spherical harmonics (SH) function, is employed to convert an appearance feature vector and a viewing direction $d$ into the final color $c$. However, in this study, we will take advantage of the variant of MLP structures to get more customization for few-shot scenarios. Those $A$ and $A_{c}$ tensor components play a pivotal role in the revolution of our Few TensorRF technique.
@@ 3.2 Enhance Few-shot on Tensorial Radiance Field (Few TensoRF)
Given the findings depicted in Figure 3, showcasing the rendered novel views with varying amounts of training images in the TensoRF baseline, it becomes evident that TensoRF encounters challenges in delivering satisfactory results under sparse input conditions. The visual comparison highlights a noticeable disparity in the quality of novel views, particularly as the number of training images decreases. As discussed in the FreeNeRF Yang et al. (2023), we anticipate encountering issues arising from over-fast convergence during the optimization process. This rapid convergence, particularly in high-frequency components, hinders TensorRF’s ability to effectively explore low frequency information. Consequently, it introduces a quite bias in TensoRF, leading to the manifestation of undesired high-frequency artifacts (TensoRF 3img in Fig. 3).
Therefore, to achieve better results in sparse input cases, the proposed Few TensoRF method proposes three improvements on the TensoRF Chen et al. (2022) baseline, as depicted in Figure 2. Frequency Masking Tensor Components. The goal here is to decrease the sensitivity of each tensor component in the high-frequency domain during the initial stages of training. This approach guides our model to concentrate on enhancing low-frequency structures, ensuring the stability of tensor components against undesired high-frequency artifacts. As mentioned earlier, there are two types of components: A for the density component and Ac for the appearance color feature component. Two frequency masks will be applied to them according to the following equations:
@@
Where $\alpha_{i}(t,T,L)$ denotes the $i^{th}$ bit value of $\alpha(t,T,L)$ ; $t$ and $T$ are the current training iteration and the total iterations, respectively, $L$ corresponds to either the density component length or the appearance feature length, depending on the desired tensor component to mask. This dynamic frequency mask, a unique feature introduced in FreeNeRF Yang et al. (2023), calculates the mask values based on the current iteration if $t$ is less than $T$. The calculation involves both an integer and a fractional part, ensuring a smooth transition. If $t$ is greater than or equal to $T$, the method returns a tensor of ones. This dynamic frequency mask is particularly beneficial for enhancing the high-frequency details of rendering novel views in the later stages of training. Additionally, we have implemented a simplified version of Eqn. 4, allowing us to set a fixed visible ratio for the frequency mask. This can be expressed as follows:
@@ 4.2 Comparison
Table 2: Training time comparison on the Hotdog scene. The bold markings indicate the most optimal results obtained in our experiments.